17. ImaChemist found the density of Freon-11 (CFCl3) to be 5.58 g/L under her experimental conditions. Hermeasurements showed that the density of an unknown gas was 4.38 g/Lunder the same conditions. What is the molar mass of theunknown?
Answer #1
m/V = 5.58 g/L = (P/RT)(137.4)
P/RT is a constant in each case......= 5.58/137.4 =0.0406
4.38 g/L = 0.0406 (unk MM)
unk MM = 4.38 / 0.0406 = 108 g/mol